# IPMU25-0011

### v1

@@ 2 Minimal $\boldsymbol{\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$ Model
Note that in a renormalizable model where there is only one complex scalar field that spontaneously breaks $\mathrm{U}(1)_{L_{\mu}-L_{\tau}}$ symmetry, the (type-I) seesaw mechanism [27, 28, *Yanagida:1979gs, 30, 31, 32] is the minimal setup to reproduce the neutrino oscillation experiments. With the gauge charges given in Tab. 1, the interaction terms responsible for neutrino mass and mixing are given by,
@@
Here, $H$ denotes the SM Higgs doublet. Due to the $\mathrm{U}(1)_{L_{\mu}-L_{\tau}}$ gauge symmetry, the Yukawa interactions of the charged leptons are diagonal, $\kappa_{e,\mu,\tau}$. Moreover, by exploiting the phase freedom of the singlet charged leptons, these Yukawa couplings can be made real and positive without loss of generality. Similarly, the Yukawa interactions $\kappa_{i}$ between the right-handed neutrinos and the lepton doublets are also diagonal, and can be rendered real and positive by utilizing the phase freedom of the lepton doublets. Additionally, the mass terms $M_{ee}$ and $M_{\mu\tau}$ can be made real and positive through the phase rotations of the right-handed neutrinos. However, it is not possible to simultaneously render both $\kappa_{e\mu}$ and $\kappa_{e\tau}$ real-valued using phase transformations. Consequently, the minimal $\mathrm{U}(1)_{L_{\mu}-L_{\tau}}$ model inherently contains a single physical parameter responsible for CP violation.
@@
The Majorana phases $\eta_{1,2}$ are fixed to ensure positive-definite neutrino masses.
@@ 3 Analysis and Result
NuFIT 6.0 - - - $2.7$ $90\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 KATRIN - - $2.9$ $90\%$ CL
NuFIT 6.0 - $M^{0\nu}=1.11$ $5.1$ $90\%$ CL
NuFIT 6.0 - $M^{0\nu}=2.39$ $7.6$ $90\%$ CL
@@
NuFIT 6.0 - - - $4.3$ $96\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 KATRIN - - $4.4$ $96\%$ CL

### v2

@@ 2 Minimal $\boldsymbol{\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$ Model
Note that in a renormalizable model where there is only one complex scalar field that spontaneously breaks $\mathrm{U}(1)_{L_{\mu}-L_{\tau}}$ symmetry, the (type-I) seesaw mechanism [27, 28, *Yanagida:1979gs, 30, 31, 32] is the minimal setup to reproduce the neutrino oscillation experiments.333 For other $\mathrm{U}(1)_{L_{\mu}-L_{\tau}}$ models based on alternative seesaw mechanisms, see the Appendix. With the gauge charges given in Tab. 1, the interaction terms responsible for neutrino mass and mixing are given by,
@@
Here, $H$ denotes the SM Higgs doublet. Due to the $\mathrm{U}(1)_{L_{\mu}-L_{\tau}}$ gauge symmetry, the Yukawa interactions of the charged leptons are diagonal, $\kappa_{e,\mu,\tau}$. Moreover, by exploiting the phase freedom of the singlet charged leptons, these Yukawa couplings can be made real and positive without loss of generality. Similarly, the Yukawa interactions $\kappa_{i}$ between the right-handed neutrinos and the lepton doublets are also diagonal, and can be rendered real and positive by utilizing the phase freedom of the lepton doublets. Additionally, the mass terms $M_{ee}$ and $M_{\mu\tau}$ can be made real and positive through the phase rotations of the right-handed neutrinos. However, it is not possible to simultaneously render both $\kappa_{e\mu}$ and $\kappa_{e\tau}$ real-valued using phase transformations. Consequently, the minimal $\mathrm{U}(1)_{L_{\mu}-L_{\tau}}$ model inherently contains a single physical parameter responsible for CP violation. We will make this point more explicit later by employing the constraint equations associated with the two-zero minor structures.
@@
The quantities $r_{21}$ and $r_{31}$ on the left-hand side of Eqs.,(2.9) and (2.10) are positive real numbers, while the right-hand side is generally complex. To ensure that the right-hand side also becomes a positive real number, two of the three CP phases $\eta_{1}$, $\eta_{2}$, and $\delta_{\mathrm{CP}}$ must be fixed in terms of the remaining one. In our analysis, we take $\delta_{\mathrm{CP}}$ as the free parameter and determine $\eta_{1}$ and $\eta_{2}$ accordingly.
@@ 3 Analysis and Result
NuFIT 6.0 - - - $4.3$ $96\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 KATRIN - - $4.4$ $96\%$ CL
NuFIT 6.0 - $M^{0\nu}=1.11$ $5.1$ $96\%$ CL
NuFIT 6.0 - $M^{0\nu}=2.39$ $7.6$ $96\%$ CL
@@
NuFIT 6.0 - - - $2.7$ $90\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 KATRIN - - $2.9$ $90\%$ CL
@@ 4 Conclusion
In this work, we have primarily focused on the most conventional realization of the seesaw mechanism, adopting it as a minimal model. Here, we briefly comment on alternative neutrino mass generation mechanisms. If neutrinos are Dirac fermions, the presence of ${\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$ symmetry enforces a diagonal mass matrix at the renormalizable level, rendering the observed neutrino mixing angles unexplained within such a framework. For the type-III seesaw mechanism [52], where fermionic SU(2) triplets are introduced to generate neutrino masses, the arguments presented in this paper apply straightforwardly, leading to the same constraints on the neutrino mass matrix. In contrast, the minimal type-II seesaw model [53, 54, 55, 56], where an SU(2) triplet scalar is responsible for generating neutrino masses, is not consistent with current experimental data under ${\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$ symmetry. However, by extending the model to include multiple SU(2) triplet scalars that transform appropriately under ${\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$, it is possible to obtain a richer structure for the neutrino mass matrix. An analysis of such extended scenarios is presented in the Appendix.
@@ Acknowledgements
Appendix A Constraints on Two-Zero Textures in $\boldsymbol{\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$ Model
In the minimal type-I seesaw model, the ${\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$ symmetry imposes a constraint on the inverse neutrino mass matrix, requiring that $m_{\nu}^{-1}$ has vanishing $(\mu,\mu)$ and $(\tau,\tau)$ entries. In contrast, in several non-minimal neutrino mass generation scenarios, the same ${\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$ symmetry can directly constrain the neutrino mass matrix itself, leading to
$(m_{\nu})_{\mu\mu}=(m_{\nu})_{\tau\tau}=0\,.$ (A.1)
Such relations can arise, for example, in type-II seesaw models where three SU(2) triplet scalars with ${\mathrm{U}(1)_{L_{\mu}-L_{\tau}}}$ charges 0, $+1$, and $-1$ are introduced to generate the neutrino mass matrix. Similar structures can also be realized in radiative neutrino mass generation mechanisms [57, 58] and in inverse seesaw models [59]. This condition again imposes relations among $m_{1}$, $m_{2}$, and $m_{3}$, allowing for statistical analyses analogous to those presented in the main text. Figure 7 shows the $\Delta\chi^{2}$ as a function of the total neutrino mass $m_{\mathrm{tot}}$ for the NO and IO cases. In the NO case, neutrino oscillation data imposes a lower bound of $m_{\mathrm{tot}}\gtrsim 0.5$ eV, which is in significant tension with cosmological constraints. For the IO case, this lower bound is relaxed to $m_{\mathrm{tot}}\gtrsim 0.2$ eV, resulting in a milder, though still non-negligible—conflict with cosmological observations, particularly with BAO data. The resulting constraints for both cases are summarized in Tables 4 and 5. The degrees of freedom used in the $\chi^{2}$ analysis for each dataset are the same as those described in the main text.
(a) NO
(b) IO
Figure 7: The values of $\Delta\chi^{2}$ incorporating neutrino oscillation data, cosmological observations, and direct neutrino mass measurements, under the condition $(m_{\nu})_{\mu\mu}=(m_{\nu})_{\tau\tau}=0$.
Table 4: Constraints on neutrino properties under the two-zero texture condition $(m_{\nu})_{\mu\mu}=(m_{\nu})_{\tau\tau}=0$ in the NO case.
Oscillation $m_{\beta}$ Cosmology $\Delta\chi^{2}$(NO) Confidence Level
IC19 w/o SK-atm (NuFIT 6.0)
NuFIT 6.0 - - $3.6$ $94\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 KATRIN - $4.1$ $94\%$ CL
NuFIT 6.0 - Planck $13$ $3.2\sigma$
NuFIT 6.0 - Planck + DESI $57$ $7.2\sigma$
IC24 with SK-atm (NuFIT 6.0)
NuFIT 6.0 - - $2.0$ $84\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 KATRIN - $2.6$ $84\%$ CL
NuFIT 6.0 - Planck $12$ $3.0\sigma$
NuFIT 6.0 - Planck + DESI $55$ $7.1\sigma$
Table 5: Same as Tab. 4, but for the IO case.
Oscillation $m_{\beta}$ Cosmology $\Delta\chi^{2}$(IO) Confidence Level
IC19 w/o SK-atm (NuFIT 6.0)
NuFIT 6.0 - - $1.4$ $76\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 KATRIN - $1.5$ $76\%$ CL
NuFIT 6.0 - Planck $2.9$ $77\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 - Planck + DESI $8.9$ $2.5\sigma$
IC24 with SK-atm (NuFIT 6.0)
NuFIT 6.0 - - $0.6$ $56\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 KATRIN - $0.7$ $56\%$ CL
NuFIT 6.0 - Planck $2.4$ $70\%$ ${}^{\phantom{\dagger}}$
NuFIT 6.0 - Planck + DESI $9.7$ $2.6\sigma$